OpenStudy (anonymous):

heeeelp please! The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2). y = −5x + 13 y = negative 1 over 5x + 7 over 5 y = 1 over 5x − 7 over 5 y = 5x − 17

OpenStudy (studygurl14):

what do you know about the slopes of perpendicular lines?

OpenStudy (anonymous):

yes ma'am

OpenStudy (anonymous):

I just don't understand slope equations

OpenStudy (anonymous):

In y = mx + b form, the m position is the slope. So look at your equation, y = 5x + 4, and tell me what the slope is

OpenStudy (anonymous):

5 is the slope

OpenStudy (anonymous):

correct....however, we are looking for a perpendicular line. We need the negative reciprocal of the slope given. All that means is " flip " the slope, and change the sign. For example...you have a slope of 1/7 and you want the negative reciprocal of that slope. Flip the slope and change the sign. The negative reciprocal of 1/7 is -7/1. So what do you think the negative reciprocal of 5 (or 5/1) is ?

OpenStudy (anonymous):

-5 or -5/1 right?

OpenStudy (anonymous):

negative reciprocal of 5 (or 5/1) is -1/5...do you see how I flipped the slope and changed the sign ?

OpenStudy (anonymous):

yes I do, thank you

OpenStudy (anonymous):

so next I would...

OpenStudy (anonymous):

we can now either use point slope formula--y - y1 = m(x - x1, or slope intercept formula--y = mx + b, to find the perpendicular equation. Which would you prefer to use

OpenStudy (anonymous):

either one will give you the same answer...so which one do you want to use ?

OpenStudy (anonymous):

hello ??

OpenStudy (anonymous):

the second one seems less complicated lol

OpenStudy (anonymous):

so --y = mx + b I guess

OpenStudy (anonymous):

-y= -5/1 + 4 ?

OpenStudy (anonymous):

true....slope intercept formula : y = mx + b slope(m) = -1/5 (-3,2) x = -3 and y = 2 we already know the slope, so basically we will be solving for b (the y intercept) Lets sub now... y = mx + b 2 = -1/5(-3) + b -- distribute through the parenthesis 2 = 3/5 + b -- subtract 3/5 from each side 2 - 3/5 = b -- common denominator is 10 20/10 - 6/10 = b 14/10 = b 7/5 = b so our equation is : y = -1/5x + 7/5 any questions ?

OpenStudy (anonymous):

wow ive never understood this until now thank you so much! do you just answer these questions in your free time bc that's really nice of you!

OpenStudy (anonymous):

I do not mind helping out....there might be a time when I need somebody to help me :)

OpenStudy (anonymous):

well thanks again :)))

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

so do I do the same thing to this problem? The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5)

OpenStudy (anonymous):

One thing you need to remember...when looking for perpendicular slope, you need the negative reciprocal of slope given. But when you need a parallel slope, the slope remains the same.

OpenStudy (anonymous):

so what is the slope of the equation given ? Thats the first thing you need to find

OpenStudy (anonymous):

remember, in y = mx + b form, the slope is in the m position. So do you know what the slope of y = -2x - 2 is ?

OpenStudy (anonymous):

well its y= -2x-2

OpenStudy (anonymous):

slope is -2

OpenStudy (anonymous):

correct....and since we are looking for a parallel line, the slope will remain the same. Now we can use y = mx + b form slope(m) = -2 (4,5) x = 4 and y = 5 now sub and solve for b. Can you do that ?

OpenStudy (anonymous):

so I make it 5=-2x(4)-2

OpenStudy (anonymous):

or -b?

OpenStudy (anonymous):

5 = -2(4) + b --- remember you subbed in 4 for x, so you do not need to list the variable x......and it is +b....because y = mx + b. Understand ?

OpenStudy (anonymous):

yes yes I do

OpenStudy (anonymous):

now can you finish and solve for b ?

OpenStudy (anonymous):

leme try brb

OpenStudy (anonymous):

does b=13

OpenStudy (anonymous):

5 = -2(4) + b 5 = -8 + b 5 + 8 = b 13 = b correct....now we know that the slope is -2....so our equation will be : y = -2x + 13 Understand ?

OpenStudy (anonymous):

yes! finally lolol thanks a million

OpenStudy (anonymous):

sure thing...glad I could help :)

OpenStudy (anonymous):

:)